# Low-temperature thermal transport in moiré superlattices

### v1

@@ Abstract
We calculate the thermal conductivity of various moiré bilayer systems using a continuum approach and the semiclassical transport theory. When the twist angle is close to 0, we observe a significant reduction of thermal conductivity in a particular low-temperature regime. This reduction is attributed to a moiré-induced reconstruction of acoustic phonon bands and associated decrease of the group velocity. Conversely, in the zero temperature limit, the thermal conductivity is enhanced by moiré effect, surpassing the original values in non-moiré counterparts. These changes result in a characteristic temperature dependence which deviates from the quadratic behavior in intrinsic two-dimensional systems.
@@ I Introduction
Heat transport in 2D materials has attracted notable interest because of its superiority and its sensitivity to heterostructuring. For example, the high thermal conductivity of suspended graphene [34, 35, 36, 37] is significantly decreased in the presence of substrate or other layer due to strong scattering of flexural phonons [38, 39, 40, 41, 42, 43, 44]. Meanwhile, experimental measurements demonstrated further suppression when rotationally stacked graphene layers [45, 46]. Theoretical simulations showed that such a thermal conductivity reduction by a twist can be caused by the enhancement of anharmonic phonon scattering and the redistribution of phonons to higher frequencies [47, 48, 49, 50, 51]. While all these simulations assumed the room temperature and above, the effect of moiré-induced phonon band reconstruction on the thermal properties is expected to be relevant in the low temperature regime, and it remains unknown.
@@
In this paper, we calculate the thermal conductivity of various representative twisted bilayer systems, including twisted bilayer graphene (TBG) and twisted graphene/hexagonal boron nitride (t-G/hBN), and also twisted bilayer molybdenum disulfide (t-MoS2) in both parallel (P) and antiparallel (AP) stacking arrangement [Fig. 1]. Here, we obtain the phonon band structure by the continuum approach, which has been extensively utilized to describe lattice relaxation and the formation of superlattice mini-bands in the phonon dispersion relation [23, 24, 52, 25]. Then we calculate the thermal conductivity using a semiclassical transport theory in the low temperature regime assuming a constant mean free path. We demonstrate that the flattening of the low-energy phonons bands leads to a significant suppression of thermal conductivity up to 40%, resulting in a characteristic deviation from the generic quadratic temperature dependence of thermal conductivity in two-dimensional system.
@@ II.1 Lattice geometry
For honeycomb lattice components, we take graphene, hexagonal boron nitride, and molybdenum disulfide with lattice constant of $a\approx$ 0.246 nm, 0.2504 nm, and 0.317 nm, respectively. In this paper, we consider twisted bilayer graphene (TBG) and twisted bilayer molybdenum disulfide (t-MoS2) as examples of homo-bilayer ($\varepsilon=1$) and also twisted graphene/hexagonal boron nitride (t-G/hBN) as a heterobilayer ($\varepsilon\neq 1$). In Fig. 2(c), we illustrate the formation of moiré superlattice in t-G/hBN with $\theta=1.25^{\circ}$ where a lattice constant difference $\varepsilon\approx 1.8\%$ and the twist angle produce a moiré pattern of $L_{\rm M}=|\mathbf{L}^{\rm M}_{i}|\approx 8.8$ nm.
Across the moiré pattern, the local stacking structure changes smoothly at the atomic scale. At a given position $\mathbf{r}$, it is characterized by the phase difference $(\varphi_{1},\varphi_{2})$ defined as
@@
For example, $(\varphi_{1},\varphi_{2})=(0,0)$, $(2\pi/3,2\pi/3)$ and $(4\pi/3,4\pi/3)$ correspond to AA (complete alignment of the honeycomb lattices), AB (B-site of layer 2 on top of A-site of layer 1), and BA (A-site of layer 2 on top of B-site of layer 1), respectively [Fig. 2(b)]. We note that, when stacking two monolayers with different sublattice atoms, e.g., t-MoS2, there are two possible stacking configuration, parallel (P) and antiparallel (AP), which are related by 180 rotation of layer 1. In the parallel stacking, the two layers have identical atoms at A and A (B and B), whereas in the antiparallel stacking, the two layers have different atoms on the same sublattice sites.
@@ II.2 Continuum model
We describe the in-plane lattice vibration in the twisted bilayer systems using a continuum approach [23, 24, 25]. In this framework, we consider a smoothly-varying displacement field $\mathbf{u}^{(l)}(\mathbf{r},t)$ at layer $l(=1,2)$, which represents the shift of the atom at $\mathbf{r}$ in the original rigid lattice, and express the Lagrangian as a functional of $\mathbf{u}^{(l)}(\mathbf{r},t)$. Here we only consider the in-plane atomic displacement $\mathbf{u}^{(l)}=(u_{x}^{(l)},u_{y}^{(l)})$ neglecting the out-of-plane component. In the presence of the lattice distortion, the phase difference [Eq. (2)] is modified as
@@
The Lagrangian is explicitly written as $L=T-(U_{E}+U_{B})$, where $T$ is the kinetic energy, $U_{E}$ is the elastic energy, and $U_{B}$ is the interlayer binding energy. The $U_{B}$ is defined by
$U_{B}=\int V[\varphi_{1}(\mathbf{r},t),\varphi_{2}(\mathbf{r},t)]\,d^{2}\mathbf{r},$ (4)
where $V[\varphi_{1},\varphi_{2}]$ is the interlayer binding energy per area of the non-rotated bilayer with an interlayer registry specified by $\varphi_{i}$. By 120 symmetry of the system, it should be written within the lowest harmonics as
$V[\varphi_{1},\varphi_{2}]=\sum_{j=1}^{3}2V_{0}\cos\left(\varphi_{j}+\varphi_{0}\right),$ (5)
where $\varphi_{3}=-\varphi_{1}-\varphi_{2}$, and the constant energy offset is neglected. The energy of AA, AB, and BA bilayer stacking is then given by $V_{AA}=6V_{0}\cos(\varphi_{0})$, $V_{AB}=6V_{0}\cos(\varphi_{0}+\tfrac{2}{3}\pi)$, and $V_{BA}=6V_{0}\cos(\varphi_{0}-\tfrac{2}{3}\pi)$. Here we obtain the parameters $(V_{0},\varphi_{0})$ from the relative values of $V_{AA}$, $V_{AB}$, and $V_{BA}$ found in the literature. Table 1 lists the values of $(V_{0},\varphi_{0})$ and the corresponding $(V_{AA},V_{AB},V_{BA})$ for the considered systems in this paper.
@@
where $\rho^{(l)}$ is the mass density of layer $l$. The mass density for graphene, hBN, and MoS2 are given in Table 2. For the layer convenience, we define averaged materials parameters
@@
II.3 Thermal transport theory
Based on linearized Boltzmann transport theory with relaxation time approximation [64], thermal conductivity of 2D isotropic material can be expressed as
@@ II.3 Thermal transport theory
where $S$ is the total area of the system, $f_{0}(\omega)=1/[\exp(\hbar\omega/k_{B}T)-1]$ is the Bose-Einstein distribution function, $v_{n,\mathbf{q}}=|\bm{\nabla}_{\mathbf{q}}\omega_{n,\mathbf{q}}|$ is the phonon velocity, $\tau_{n,\mathbf{q}}$ is the relaxation time, and the summation is taken over all of the mode index $n$ and wave vector $\mathbf{q}$. Here, we note that the isotropic behavior is a consequence of the three-fold rotation symmetry.
@@
where $\beta=1/(k_{B}T)$. In Eq. (19), $\tilde{n}(\omega)$ contains all information regarding the phonon dispersion while $C(\omega,T)$ acts as a weight function. The function $C(\omega,T)$ is plotted in Fig. 4; it equals to $k_{B}$ when $\omega=0$, and decays exponentially for $\hbar\omega\gtrsim 2k_{B}T$. In general, the summation over mode index $n$ would include both in-plane and out-of-plane (flexural) phonon modes. However, scattering due to substrate or other layers significantly reduces the mean free path of the flexural mode [39, 40, 41, 42, 43, 44], hence, its contribution to thermal conductivity especially at low temperatures [38]. Hereafter, we will focus only on the in-plane modes contribution in the thermal conductivity.
Figure 4: Plot of the spectral heat capacity function $C(\omega,T)$ [Eq. (21)] as a function of phonon energy at different temperatures.
Since only the antisymmetric phonon modes are strongly affected by the interlayer moiré potential, it is useful to write $\kappa$ and $\tilde{n}$ in terms of the symmetric and antisymmetric components
@@
where ${\bar{v}}^{-1}=v_{\rm L}^{-1}+v_{\rm T}^{-1}$, $\zeta(n)$ is the Riemann zeta function, and NC stands for ‘non-coupled’. Here, we note that we have neglected the out-of-plane (flexural) phonon modes in the calculation of thermal conductivity. The quadratic temperature dependence of the thermal conductivity is a characteristic of linear acoustic phonon-dominated thermal transport of 2D systems in the low-temperature regime [38].
@@ III.1 TBG and t-G/hBN
Figure 5 shows the calculated band dispersion and density of states for antisymmetric phonon modes in TBG with $\theta=$ 5, 2.65 and 0.817. Figure 6 presents similar plots for t-G/hBN with $\theta=$ 5, 3, and 0. The phonon dispersion is plotted along the high-symmetry line in the moiré Brillouin zone (MBZ) as illustrated in Fig. 2(c). In each panel, the red-dashed line indicate the dispersion (only the lowest two branches shown) and the density of states in the non-coupled case. Here we can see that the interlayer coupling significantly modifies the band structure of the antisymmetric modes in the energy range below $\sim$20 meV. This reconstruction is characterized by sharp peaks in the density of states which results from the flattening of superlattice phonon bands [23, 24, 25, 32]. As the twist angle is reduced towards 0, the number of sharp peaks increases dramatically. On the other hand, the phonon bands of the symmetric phonon modes are not affected by the moiré interlayer coupling, hence they are equivalent to those of the non-coupled case (red-dashed lines).
@@
and the color brightness is a linear scale of $D(\omega,v)$. The horizontal red-dashed lines correspond to the velocity of the TA ($v_{\rm T}$) and LA ($v_{\rm L}$) phonons of a single layer [Eq. (9)]. Here, the flattening of phonon bands can be immediately seen as a distribution of phonon modes below $v_{\rm T}$ and $v_{\rm L}$. In 0.817 TBG [Fig. 5(f)] and 0 t-G/hBN [Fig. 6(f)], particularly, the signals in $\hbar\omega<15$ meV predominantly falls below the line of $v_{\rm T}$, i.e., nearly all of the low-energy antisymmetric phonons become slower than the original acoustic phonons in its non-moiré counterpart. Note that the distribution of symmetric phonon states (not shown) sticks to the $v_{L}$ and $v_{T}$ lines, where the intensity increases linearly with energy.
@@
Figure 7: (a) Velocity-weighted density of states (VDOS) of the antisymmetric modes, $\tilde{n}^{-}(\omega)$, in TBGs with various twist angles. Black dashed lines represents $\tilde{n}^{+}_{\rm NC}=\tilde{n}^{-}_{\rm NC}$ for a non-coupled bilayer [Eq. (23)]. The VDOS for symmetric phonons $\tilde{n}^{+}$ is equal to $\tilde{n}^{\pm}_{\rm NC}$. (b) Similar plots for t-G/hBN.
@@
Figure 7(a) and (b) show $\tilde{n}^{-}(\omega)$ (the VDOS of the antisymmetric phonon modes) of TBG and t-G/hBN, respectively, with various small twist angles. The effect of the moiré coupling is observed as a difference from a black-dashed line, which represents $\tilde{n}^{+}_{\rm NC}=\tilde{n}^{-}_{\rm NC}\propto\omega$ [Eq. (23)]. While VDOS is proportional to both phonon density of states and velocity, we find that the reduction of phonon velocity is more significant than the sharpening of density of states, leading to a suppression of $\tilde{n}^{-}$ over a wide range of phonon frequency. We also find that the linear behavior of $\tilde{n}^{-}$ remains in the low frequency region, which corresponds to the linear dispersion in the the lowest moiré phonon band in $\omega<\omega_{\rm edge}\approx 2\pi v_{0}/L_{\rm M}$. For the symmetric phonon modes (not shown), we have $\tilde{n}^{+}(\omega)=\tilde{n}^{\pm}_{\rm NC}(\omega)$, because they are not influenced by the moiré coupling.
@@
Figure 8(a) and (b) summarise the calculated thermal conductivity $\kappa(T)(=\kappa^{+}+\kappa^{-})$ of TBG and t-G/hBN, respectively. In each figure, the left panel shows a log-log plot of $\kappa(T)$ for $1002<T<$ 100 K. Here the vertical axis is scaled by the constant mean-free-path length $\Lambda$, which is assumed to be independent of temperature. The colored lines correspond to different twist angles. The black-dashed line represents the non-coupled bilayer case, $\kappa_{\rm NC}$, which is proportional to $T^{2}$ as given in Eq. (24). Here, we find that thermal conductivity is suppressed from the non-coupled case, notably around 20 K, while it converges towards the non-coupled value as increasing temperatures.
@@
To better understand the change from the intrinsic graphene, we plot the relative thermal conductivity, $\kappa(T)/\kappa_{\rm NC}(T)$, in the right panel of Fig. 8(a) and (b). Here, suppression of thermal conductivity is represented by a value below unity. We can see that the suppression occurs over a wide temperature range except near the low-temperature limit. This suppression becomes more pronounced as the twist angle is smaller, where the largest reduction of up to $\sim$35for 0.817 TBG and up to $\sim$40for 0 t-G/hBN takes place at around 20 K. This can be understood since at around 20 K, heat transport is carried by phonons with energy below $\sim$20 meV [see Fig. 4], where the moiré effect on the asymmetric modes are the most notable, as shown by the changes in VDOS [Fig. 7]. There, thermal conductivity is almost entirely came from the symmetric phonon modes, $\kappa^{+}=\kappa_{\rm NC}/2$ (red-dashed line).
@@
The calculated phonon dispersion for t-MoS2 is shown in Fig. 9 for both parallel and antiparallel stacking case with twist angle 0.8. We find that the antisymmetric phonon modes of the parallel-stacked case closely resembles that of TBG [Fig. 5(c)]. This similarity occurs because TBG and t-MoS2 share triangular domain wall structures [see Fig. 3(a) and (c)]. In an analogous way, the phonon band structure of the antiparallel t-MoS2 resembles those of t-G/hBN, reflecting a common honeycomb domain-wall structure [Fig. 3(b) and (d)]. This is a natural result because the moiré phonon band structure is qualitatively reproduced by an effective mass-spring model for domain-wall motion [25], and hence it is primarily determined by the geometical structure of domain walls (triangular or honeycomb). We also observe that the characteristic energy scale of the moiré phonon bands in t-MoS2 is much smaller than in TBG and t-G/hBN, because the original acoustic phonon velocity in MoS2 is much lower than that of graphene and hBN.

### v2

@@ Abstract
We calculate the phonon thermal conductivity of various moiré bilayer systems using a continuum approach and the semiclassical transport theory. When the twist angle is close to 0, we observe a significant reduction of thermal conductivity in a particular low-temperature regime. This reduction is attributed to a moiré-induced reconstruction of acoustic phonon bands and associated decrease of the group velocity. Conversely, in the zero temperature limit, the thermal conductivity is enhanced by moiré effect, surpassing the original values in non-moiré counterparts. These changes result in a characteristic temperature dependence which deviates from the quadratic behavior in intrinsic two-dimensional systems.
@@ I Introduction
Heat transport in 2D materials has attracted notable interest because of its superiority and its sensitivity to heterostructuring. For example, the high thermal conductivity of suspended graphene [34, 35, 36, 37] is significantly decreased in the presence of substrate or other layer due to strong scattering of flexural phonons [38, 39, 40, 41, 42, 43, 44]. Meanwhile, experimental measurements demonstrated further suppression in rotationally stacked graphene layers [45, 46]. Theoretical simulations suggested that such a thermal conductivity reduction by a twist occurs due to the enhancement of anharmonic phonon scattering or the redistribution of phonons towards higher frequencies [47, 48, 49, 50, 51]. However, these studies have focused solely on room temperature and above, while the effect of moiré-induced phonon band reconstruction is expected to be relevant in the low-temperature regime, and it remains unknown.
@@
In this paper, we calculate the phonon thermal conductivity of various representative twisted bilayer systems, including twisted bilayer graphene (TBG) and twisted graphene/hexagonal boron nitride (t-G/hBN), and also twisted bilayer molybdenum disulfide (t-MoS2) in both parallel (P) and antiparallel (AP) stacking arrangement [Fig. 1]. Here, we obtain the phonon band structure by the continuum approach, which has been extensively utilized to describe lattice relaxation and the formation of superlattice mini-bands in the phonon dispersion relation [23, 24, 52, 25]. Then we calculate the thermal conductivity using a semiclassical transport theory in the low temperature regime assuming a constant mean free path. We demonstrate that the flattening of the low-energy phonons bands leads to a significant suppression of thermal conductivity up to 40%, resulting in a characteristic deviation from the generic quadratic temperature dependence of thermal conductivity in two-dimensional system.
@@ II.1 Lattice geometry
For honeycomb lattice components, we take graphene, hexagonal boron nitride, and molybdenum disulfide with lattice constant of $a\approx$ 0.246 nm, 0.2504 nm, and 0.317 nm, respectively. In this paper, we consider twisted bilayer graphene (TBG) and twisted bilayer molybdenum disulfide (t-MoS2) as examples of homo-bilayer ($\varepsilon=0$) and also twisted graphene/hexagonal boron nitride (t-G/hBN) as a heterobilayer ($\varepsilon\neq 0$). In Fig. 2(c), we illustrate the formation of moiré superlattice in t-G/hBN with $\theta=1.25^{\circ}$ where a lattice constant difference $\varepsilon\approx 1.8\%$ and the twist angle produce a moiré pattern of $L_{\rm M}=|\mathbf{L}^{\rm M}_{i}|\approx 8.8$ nm.
Across the moiré pattern, the local stacking structure changes smoothly at the atomic scale. At a given position $\mathbf{r}$, it is characterized by the phase difference
$\bm{\varphi}(\mathbf{r})=(\varphi_{1}(\mathbf{r}),\varphi_{2}(\mathbf{r}))$, where $\varphi_{j}(\mathbf{r})$ is defined as
@@
For example, $\bm{\varphi}=(0,0)$, $(2\pi/3,2\pi/3)$ and $(4\pi/3,4\pi/3)$ correspond to AA (complete alignment of the honeycomb lattices), AB (B-site of layer 2 on top of A-site of layer 1), and BA (A-site of layer 2 on top of B-site of layer 1), respectively [Fig. 2(b)].
We note that, when stacking two monolayers with different sublattice atoms, e.g., t-MoS2, there are two possible stacking configuration, parallel (P) and antiparallel (AP), which are related by 180 rotation of layer 1. In the parallel stacking, the two layers have identical atoms at A and A (B and B), whereas in the antiparallel stacking, the two layers have different atoms on the same sublattice sites.
@@ II.2 Continuum model
We describe the in-plane lattice vibration in the twisted bilayer systems using a continuum approach [23, 24, 25]. In this framework, we consider a smoothly-varying displacement field $\mathbf{u}^{(l)}(\mathbf{r},t)$ at layer $l(=1,2)$, which represents the shift of the atom at $\mathbf{r}$ in the original rigid lattice, and express the Lagrangian as a functional of $\mathbf{u}^{(l)}(\mathbf{r},t)$. Let us consider the in-plane atomic displacement $\mathbf{u}^{(l)}=(u_{x}^{(l)},u_{y}^{(l)})$. In the presence of the lattice distortion, the phase difference [Eq. (2)] is modified as
@@
The Lagrangian is explicitly written as $L=T-(U_{E}+U_{B})$, where $T$ is the kinetic energy, $U_{E}$ is the elastic energy, and $U_{B}$ is the interlayer binding energy. The total interlayer binding energy is calculated as the integral of the local interlayer binding energy over the whole system,
$U_{B}=\int V[\bm{\varphi}(\mathbf{r},t)]\,d^{2}\mathbf{r},$ (4)
where $V[\bm{\varphi}]$ is the interlayer binding energy per area of the non-rotated bilayer with an interlayer registry specified by $\bm{\varphi}$ [53, 25]. By 120 symmetry of the system, $V[\bm{\varphi}]$ should be written within the lowest harmonics as
$V[\bm{\varphi}(\mathbf{r},t)]=\sum_{j=1}^{3}2V_{0}\cos\left[\varphi_{j}\left(\mathbf{r},t\right)+\varphi_{0}\right],$ (5)
where $\varphi_{3}=-\varphi_{1}-\varphi_{2}$, $\varphi_{0}$ is a material-dependent phase, and the constant energy offset is neglected. The energy of AA, AB, and BA bilayer stacking is then given by $V_{AA}=6V_{0}\cos(\varphi_{0})$, $V_{AB}=6V_{0}\cos(\varphi_{0}+\tfrac{2}{3}\pi)$, and $V_{BA}=6V_{0}\cos(\varphi_{0}-\tfrac{2}{3}\pi)$. Here we obtain the parameters $(V_{0},\varphi_{0})$ from the relative values of $V_{AA}$, $V_{AB}$, and $V_{BA}$ found in the literature. Table 1 lists the values of $(V_{0},\varphi_{0})$ and the corresponding $(V_{AA},V_{AB},V_{BA})$ for the considered systems in this paper.
@@
where $\rho^{(l)}$ is the mass density of layer $l$. The mass density for graphene, hBN, and MoS2 are given in Table 2. For convenience, we define averaged materials parameters
@@
II.3 Phonon thermal conductivity
Based on linearized Boltzmann transport theory with relaxation time approximation [64], the phonon thermal conductivity of 2D isotropic material can be expressed as
@@ II.3 Phonon thermal conductivity
where $S$ is the total area of the system, $f_{0}(\omega)=1/[\exp(\hbar\omega/k_{B}T)-1]$ is the Bose-Einstein distribution function, $v_{n,\mathbf{q}}=|\bm{\nabla}_{\mathbf{q}}\omega_{n,\mathbf{q}}|$ is the phonon velocity, $\tau_{n,\mathbf{q}}$ is the relaxation time, and the summation is taken over all of the mode index $n$ and wave vector $\mathbf{q}$. Here, we note that the isotropic behavior is a consequence of the three-fold rotation symmetry and the thickness of the 2D system had been factored out from the expression of $\kappa$.
@@
where $\beta=1/(k_{B}T)$. In Eq. (19), $\tilde{n}(\omega)$ contains all information regarding the phonon dispersion while $C(\omega,T)$ acts as a weight function. The function $C(\omega,T)$ is plotted in Fig. 4; it equals to $k_{B}$ when $\omega=0$, and decays exponentially for $\hbar\omega\gtrsim 2k_{B}T$.
In general, the thermal conductivity is contributed by all phonon modes including the out-of-plane (flexural) phonons as well as in-plane modes. In this work, we will treat in-plane and out-of-plane phonon modes independently, since the coupling between these only appears as higher-order terms of the equation of motion [32], leading to only minor effects on the overall phonon spectrum. Also, the effect of the moiré pattern on out-of-plane phonons is expected to be smaller than that on in-plane phonons, since interlayer in-plane displacement directly modifies the local atomic registry in the moiré pattern, while the out-of-plane displacement does not [23]. Thus, we will focus on the effect of in-plane phonons on the thermal conductivity, assuming that the contribution of out-of-plane phonons remains at the intrinsic value for a single layer. It is also known that in multilayered systems or in the presence of substrate, the out-of-plane modes are strongly suppressed [40, 41, 42, 43, 44] especially at low temperature [38].
Figure 4: Plot of the spectral heat capacity function $C(\omega,T)$ [Eq. (21)] as a function of $\beta\hbar\omega$ where $\beta=1/(k_{B}T)$.
Since only the in-plane antisymmetric phonon modes are strongly affected by the interlayer moiré potential, it is useful to write $\kappa$ and $\tilde{n}$ in terms of the symmetric and antisymmetric components
@@
where ${\bar{v}}^{-1}=v_{\rm L}^{-1}+v_{\rm T}^{-1}$, $\zeta(n)$ is the Riemann zeta function, and NC stands for ‘non-coupled’. The quadratic temperature dependence of the thermal conductivity is a characteristic of linear acoustic phonon-dominated thermal transport of 2D systems in the low-temperature regime [38].
@@ III.1 TBG and t-G/hBN
Figure 5 shows the calculated band dispersion and density of states for antisymmetric phonon modes in TBG with $\theta=$ 5, 2.65 and 0.817. Figure 6 presents similar plots for t-G/hBN with $\theta=$ 5, 3, and 0. The phonon dispersion is plotted along the high-symmetry line in the moiré Brillouin zone (MBZ) as illustrated in Fig. 2(c). In each panel, the red-dashed line indicate the dispersion (only the lowest two branches shown) and the density of states in the non-coupled case. Here we can see that the interlayer coupling significantly modifies the band structure of the antisymmetric modes in the energy range below $\sim$20 meV. This reconstruction is characterized by sharp peaks in the density of states due to the band flattening, which becomes more significant at smaller twist angle. At small angle, in particular, mini band gaps also emerge at the few-meV scale. These gaps are attributed to the frequency modulations of string-like excitations that dominate the low-frequency modes of moiré superlattice systems [25]. Meanwhile, since the moiré interlayer potential does not couple the symmetric modes, the corresponding band structure is equivalent to those of the non-coupled case (red-dashed lines).
@@
Here, the Dirac delta function is approximated by a Lorentzian function where the full-width at half-maximum is between 0.02 and 0.05 meV depending on the twist angle.
@@
and the color brightness is a linear scale of $D(\omega,v)$. The horizontal red-dashed lines correspond to the velocity of the TA ($v_{\rm T}$) and LA ($v_{\rm L}$) phonons of a single layer [Eq. (9)]. The flattening of phonon bands can be immediately seen as a distribution of phonon modes below $v_{\rm T}$ and $v_{\rm L}$. In 0.817 TBG [Fig. 5(f)] and 0 t-G/hBN [Fig. 6(f)], particularly, the signals in $\hbar\omega<15$ meV predominantly falls below the line of $v_{\rm T}$, i.e., nearly all of the low-energy antisymmetric phonons become slower than the original acoustic phonons in its non-moiré counterpart. Note that the distribution of symmetric phonon states (not shown) sticks to the $v_{L}$ and $v_{T}$ lines, where the intensity increases linearly with energy.
@@
Figure 7: (a) Velocity-weighted density of states (VDOS) of the antisymmetric modes, $\tilde{n}^{-}(\omega)$, in TBGs (colored lines). Each color represent different twist angles. Colored vertical lines correspond to $\hbar\omega_{\rm edge}\approx 2\pi\hbar v_{1}/(\sqrt{3}L_{\rm M})$ for each twist angle. Black dashed lines represents $\tilde{n}^{+}_{\rm NC}=\tilde{n}^{-}_{\rm NC}$ for a non-coupled bilayer [Eq. (23)]. The VDOS for symmetric phonons $\tilde{n}^{+}$ is equal to $\tilde{n}^{\pm}_{\rm NC}$. (b) Similar plots for t-G/hBN.
@@
Figure 7(a) and (b) show $\tilde{n}^{-}(\omega)$ (the VDOS of the antisymmetric phonon modes) of TBG and t-G/hBN, respectively, with various small twist angles. The effect of the moiré coupling is observed as a difference from a black-dashed line, which represents $\tilde{n}^{+}_{\rm NC}=\tilde{n}^{-}_{\rm NC}\propto\omega$ [Eq. (23)]. While VDOS is proportional to both phonon density of states and velocity, we find that the reduction of phonon velocity is more significant than the sharpening of density of states, leading to a suppression of $\tilde{n}^{-}$ over a wide range of phonon frequency. We also find that the linear behavior of $\tilde{n}^{-}$ remains in the low frequency region, which corresponds to the linear dispersion part of the lowest moiré phonon band in $\omega<\omega_{\rm edge}\approx 2\pi v_{1}/(\sqrt{3}L_{\rm M})$ where $v_{1}$ is the phonon velocity for the lowest band near the MBZ center. For the symmetric phonon modes (not shown), we have $\tilde{n}^{+}(\omega)=\tilde{n}^{\pm}_{\rm NC}(\omega)$, because they are not influenced by the moiré coupling.
@@
Figure 8(a) and (b) summarise the calculated thermal conductivity $\kappa(T)(=\kappa^{+}+\kappa^{-})$ of TBG and t-G/hBN, respectively. In each figure, the left panel shows a log-log plot of $\kappa(T)$ for $1102<T<$ 110 K. Here the vertical axis is scaled by the constant mean-free-path length $\Lambda$, which is assumed to be independent of temperature. The colored lines correspond to different twist angles. The black-dashed line represents the non-coupled bilayer case, $\kappa_{\rm NC}$, which is proportional to $T^{2}$ as given in Eq. (24). Here, we find that thermal conductivity is suppressed from the non-coupled case, notably around 20 K, while it converges towards the non-coupled value as increasing temperatures.
@@
To better understand the change from the intrinsic graphene, we plot the relative thermal conductivity, $\kappa(T)/\kappa_{\rm NC}(T)$, in the right panel of Fig. 8(a) and (b). Here, suppression of thermal conductivity is represented by a value below unity. We can see that the suppression occurs over a wide temperature range except near the low-temperature limit. This suppression becomes more pronounced as the twist angle is smaller, where the largest reduction of up to $\sim$35for 0.817 TBG and up to $\sim$40for 0 t-G/hBN takes place at around 20 K. This can be understood since at around 20 K, heat transport is carried by phonons with energy below $\sim$20 meV ($\beta\hbar\omega\approx 6$) [see Fig. 4]. Within this energy range, the contribution of the asymmetric modes are almost completely absent [Fig. 7] and the thermal conductivity is almost entirely came from the symmetric phonon modes, $\kappa^{+}=\kappa_{\rm NC}/2$ (red-dashed line).
@@
The calculated phonon dispersion for t-MoS2 is shown in Fig. 9 for both parallel and antiparallel stacking case with twist angle 0.8. We find that the antisymmetric phonon modes of the parallel-stacked case closely resembles that of TBG [Fig. 5(c)]. This similarity occurs because TBG and t-MoS2 share triangular domain wall structures [see Fig. 3(a) and (c)]. In an analogous way, the phonon band structure of the antiparallel t-MoS2 resembles those of t-G/hBN, reflecting a common honeycomb domain-wall structure [Fig. 3(b) and (d)]. This is a natural result because the moiré phonon band structure is qualitatively reproduced by an effective mass-spring model for domain-wall motion [25], and hence it is primarily determined by the geometrical structure of domain walls (triangular or honeycomb). We also observe that the characteristic energy scale of the moiré phonon bands in t-MoS2 is much smaller than in TBG and t-G/hBN, because the original acoustic phonon velocity in MoS2 is much lower than that of graphene and hBN.